3d^2y/dx^2+dy/dx-14y=0 ; y(0)=1 , y'(0)=-1


1
Expert's answer
2022-01-30T16:25:40-0500

Characteristic (auxiliary) equation


"3r^2+r-14=0"

"D=(1)^2-4(3)(-14)=169"


"r=\dfrac{-1\pm13}{2(3)}"

"r_1=-\dfrac{7}{3}, r_2=2"

The general solution of the differential equation is


"y=c_1e^{-7x\/3}+c_2e^{2x}"

Given "y(0)=1"


"1=c_1+c_2"

"y'=-\dfrac{7}{3}c_1e^{-7x/3}+2c_2e^{2x}"

Given "y'(0)=-1"


"-\dfrac{7}{3}c_1+2c_2=-1"

"c_1=\dfrac{9}{13}, c_2=\dfrac{4}{13}"

The solution of given IVP is


"y=\dfrac{9}{13}e^{-7x/3}+\dfrac{4}{13}e^{2x}"